OpenStudy (anonymous):

Using complete sentences, explain how to graph y = one thirdx + 4 using the slope-intercept method.

OpenStudy (anonymous):

y = x + 4 Compare that to y = mx + b b = 4 and m = The slope m always has three things: 1. A sign, either + or - 2. A numerator 3. A denominator (If the slope is a whole number is has denominator 1). After one plots (0,4) then use the slope by rising 1 and running three to the right, this will land us at (3, 4). Now we connect the two points. Simple enough. :)

OpenStudy (anonymous):

Thanks a lot!! :)